Ugh!! I can not get the right answer for these....I am missing something. Can someone explain these to me?
Divide:

a. 40 X^8 Y^6 / 5 XY^3


b. X^3 –2X^2 –33X –7 / X –7


Answer by hulamc723(15)

a. 40 X^8 Y^6 / 5 XY^3
a. What we need to do in this case is look at all of the different parts of the equation separately
40/5,x^8/x, and y^6/y^3
Now simplify
40/5=8
x^8/x=x^7 (*remember x^a/x^b=x^a-b)
y^6/y^3=y^3
so now put them together again!
8x^7y^3
now im not 100% sure on the second equation and it is quite hard to explain how to figure it out online so, if i can i will but now i cant
-soo sorry